5

.

When you take a bath, how many kilograms of cold water (10°C) must you mix with hot water

(50°C) so that the temperature of the bath is 35°C? The total mass of water is 180 kg. Ignore any

heat flow between the water and its surroundings.



1
Expert's answer
2022-05-04T17:38:40-0400

"m(35-10)=(180-m)(50-35),"

"25m=180\cdot 15-15m,"

"m=67.5~kg."